You are involved in the evaluation of viewership of television shows for your local TV station.
You have been tasked to determine a 95% confidence interval for the viewing of a show called 'CSI Phoenix'.
You have randomly queried 400 local adults and determined that 150 of them watched the last episode of the show.
>> create a 95% confidence interval for the proportion of folks in the population that viewed the show
: You are involved in the evaluation of viewership of television shows for your local TV station.
You have been tasked to determine a 95% confidence interval for the viewing of a show called 'CSI Phoenix'.
You have randomly queried 400 local adults and determined that 150 of them watched the last episode of the show.
>> create a 95% confidence interval for the proportion of folks in the population that viewed the show

Answer by stanbon(18998)

You have been tasked to determine a 95% confidence interval for the viewing of a show called 'CSI Phoenix'.
----------------------------------------
You have randomly queried 400 local adults and determined that 150 of them watched the last episode of the show.
sample proportion = 150/400 = 0.375
-----------------------------------------
>> create a 95% confidence interval for the proportion of folks in the population that viewed the show
-------------------------
E = 1.96*sqrt[0.375*0.625/400] = 0.001148
-------------------------
95% CI: 0.375-0.001148 < p < 0.375+0.001148
------------------------------------------------
Cheers,
Stan H.